OpenStudy (anonymous):

What's the answer for (13x-7)=12

OpenStudy (anonymous):

Add 7 both to (13x - 7) and to 12, you obtain 13x - 7 + 7 = 12 + 7 Then 13x = 19, now divide for 13 and obtain x = 19/13 Sei italiano, Sergio? :D